If a multiple-choice test consists of 5 questions,



each with 4 possible answers of which only 1 is correct,



(a) in how many different ways can a student check off



one answer to each question?



(b) in how many ways can a student check off one



answer to each question and get all the answers



wrong?

1
Expert's answer
2022-04-10T16:20:54-0400

(a) The number of ways a student can check off each question is 4. The total number of ways he can check off all the questions:

"N=4\cdot4\cdot4\cdot4\cdot4=4^5=1024."


(b) The number of ways a student can give an incorrect answer to each question is 3. The total number of ways he can check off one answer to each question and get all the answers wrong:

"M=3^5=243."